40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 17003     Entry time: Thu Jul 14 19:09:51 2022
Author: rana 
Type: Update 
Category: General 
Subject: EQ recovery 

There was a EQ in Ridgecrest (approximately 200 km north of Caltech). It was around 6:20 PM local time.

All the suspensions tripped. I have recovered them (after some struggle with the weird profusion of multiple conflicting scripts/ directories that have appeared in the recent past...)

ETMY is still giving me some trouble. Maybe because of the HUGE bias on that within the fast CDS system, it had some trouble damping. Also the 'reenable watchdog' script in one of the many scripts directories seems to do a bad job. It re-enables optics, btu doesn't make sure that the beams are on the optical lever QPD, and so the OL servo can smash the optic around. This is not good.

Also what's up with the bashrc.d/ in some workstations and not others? Was there something wrong with the .bashrc files we had for the past 15 years? I will revert them unless someone puts in an elog with some justification for this "upgrade".

This new SUS screen is coming along well, but some of the fields are white. Are they omitted or is there something non-functional in the CDS? Also, the PD variances should not be in the line between the servo outputs and the coil. It may mislead people into thinking that the variances are of the coils. Instead, they should be placed elsewhere as we had it in the old screens.

Attachment 1: ETMY-screen.png  44 kB  Uploaded Thu Jul 14 21:16:20 2022  | Hide | Hide all
ETMY-screen.png
ELOG V3.1.3-